OpenStudy (anonymous):

Function help? f(x)=3x^2-3x and g(x)=-4x-1 find f(-2) and g(5)

OpenStudy (anonymous):

For f(-2), you substitute -2 to f(x) f(-2)=3(-2)^2 -3(-2) = 12+6 = 18 For g(5), substitute 5 to g(x) g(5)=-4(5)-1=-21

OpenStudy (anonymous):

oh, that's right... i substitute it into ALL the x's, not just the beginning. thanks!